a father is three time as old as his son in 13 years time. the father will only be twice as old as his son. find their present ages ?
Answer by ptaylor(2198)

Let x=son's present age; x+13=son's age in 13 years
Then 3x=father's present age; 3x+13=father's age in 13 years
Now we are told the following:
3x+13=2(x+13) get rid of parens
3x+13=2x+26 subtract 13 and also 2x from each side
3x-2x+13-13=2x-2x+26-13 collect like terms
x=13----------------------------son's present age
3x=3*13=39-------------------------father's present age
39+13=2*(13+13)
52=2*26
52=52
Doess this help?----ptaylor